Find the slope and y-intercept for each line
3x+5y+10=0

Answer by sprolden(40)

Find the slope and y-intercept for each line
3x+5y+10=0
1- Begin solving for y. Subtract 10 from both sides.
3x+5y+10-10=0-10
3x+5y=-10
2- Subtract 3x from both sides.
3x-3x+5y=-10-3x
5y=-10-3x
3- Divide each side by 5.
5y/5 = -10/5 - 3x/5
y = -2-3/5x
4- The solution is y=-2-3/5x OR y=-0.6x-2
5- The slope is -0.6. The y-intercept is -2.